The displacement of a particle is represented by the following equation: s = 3t3 +7t2 +5t+8 where s is in metre and t in second. The acceleration of the particle at t = 1s is


        (a)    14 m/s2


(b)    18 m/s2


        (c)    32 m/s2


        (d)    Zero


Grade:10

2 Answers

badmaash007
10 years ago

32

Jitender Pal
10 years ago

        (c)

Displacement s = 3t3 + 7t2 + 5t +8;

Velocity =  = 9t2 + 14t + 5

Acceleration =  = 18t + 14

Acceleration at

t = 1s, = 18 × 1+14 = 18 + 14 = 32m/s2